# arXiv:2601.01248v3[math.OC] 02 Mar 2026

### v3

@@ 4 Conclusion
For optimization over $\mathbb{R}^{d}$, we establish convergence of the regularized control value to the global minimum at rate $\varepsilon\ln(1/\varepsilon)$ as $\varepsilon\to 0$. For optimization over Wasserstein space $\mathcal{P}_{2}(\mathbb{R}^{d})$, we introduce a regularized mean-field control formulation and approximate it by controlled $N$-particle systems. We prove quantitative convergence to the global optimum with error of order $\frac{C_{\varepsilon}}{N}+\varepsilon\ln\!\left(\frac{1}{\varepsilon}\right),$ as $N\to\infty$ and $\varepsilon\to 0$.
@@
Under the assumptions (a) (Regularity) and (b) (Convexity) in Proposition 3.2, the stochastic Pontryagin maximum principle with both necessary and sufficient conditions established in [11, Theorems 4.5, 4.6 & 5.1] implies that, for each $(t,\mu)\in[0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d})$, $\xi\sim\mu$, the optimal control $\theta^{*}$ exists uniquely, and the optimal state process $X^{t,\xi;\theta^{*}}$ satisfies uniquely the following FBSDE:
$\begin{cases}dX^{t,\xi;\theta^{*}}_{s}=-\frac{1}{\varepsilon}Y^{t,\xi;\theta^{*}}_{s}\,ds+dW_{s},\quad X^{t,\xi;\theta^{*}}_{t}=\xi\sim\mu,\\ dY^{t,\xi;\theta^{*}}_{s}=Z^{t,\xi;\theta^{*}}_{s}\,dW_{s},\quad Y^{t,\xi;\theta^{*}}_{1}=\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1}),\\ \theta^{*}_{s}=-\frac{1}{\varepsilon}Y^{t,\xi;\theta^{*}}_{s},\quad\forall\,s\in[t,T],\end{cases}$ (A.1)
which, combined with the theories in [16, Section 5.3 and Theorem 5.7], implies that the value function $\mathcal{V}_{\varepsilon}$ is a classical solution to the master equation (1.20) satisfying Assumption 3.1 and that there holds the decoupling relation $Y^{t,\xi;\theta^{*}}_{s}=\partial_{\mu}\mathcal{V}_{\varepsilon}(s,\mathcal{L}(X^{t,\xi;\theta^{*}}_{s}))(X^{t,\xi;\theta^{*}}_{s})$ for all $s\in[t,T]$. By Lemma 3.1, the classical solution is unique.
@@
Applying Itô’s formula to the scalar product $\langle\partial_{x}X^{t,x;\theta^{*}}_{s},\partial_{x}Y_{s}^{t,x;\theta^{*}}\rangle$ and taking conditional expectation, we obtain
$\langle\partial_{x}X^{t,x;\theta^{*}}_{s},\partial_{x}Y_{s}^{t,x;\theta^{*}}\rangle$ $=\mathbb{E}\Big[\underbrace{\langle\partial_{x}X_{1}^{t,x;\theta^{*}},\,\partial_{x}\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,x,\xi;\theta^{*}}_{1})\partial_{x}X_{1}^{t,x;\theta^{*}}\rangle}_{\geq 0\text{ by Coercivity}}+\frac{1}{\varepsilon}\int_{s}^{1}|\partial_{x}Y_{r}^{t,x;\theta^{*}}|^{2}\,dr\big|\mathcal{F}_{s}\Big],$
@@
$d(|\partial_{x}X^{t,x;\theta^{*}}_{s}|^{2})=2\langle\partial_{x}X_{s}^{t,x;\theta^{*}},d(\partial_{x}X_{s}^{t,x;\theta^{*}})\rangle=-\frac{2}{\varepsilon}\langle\partial_{x}X_{s}^{t,x;\theta^{*}},\partial_{x}Y_{s}^{t,x;\theta^{*}}\rangle\,ds\leq 0.$ (A.3)
Thus, it holds that $|\partial_{x}X_{s}^{t,x;\theta^{*}}|\leq|I_{d}|=\sqrt{d}$ for all $s\in[t,1]$. Looking at $\partial_{x}Y^{t,x;\theta^{*}}$, we have
$|\partial_{x}\partial_{\mu}\mathcal{V}_{\varepsilon}(t,\mu)(x)|=|\partial_{x}Y_{t}^{t,x;\theta^{*}}|$ $=\left|\mathbb{E}\left[\partial_{x}(\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1}))\cdot\partial_{x}X^{t,x;\theta^{*}}_{1}\big|\mathcal{F}_{t}\right]\right|$
$\leq K_{G}\mathbb{E}[|\partial_{x}X^{t,x;\theta^{*}}_{1}||\mathcal{F}_{t}]$
$\leq K_{G}\sqrt{d}.$
Therefore, we obtain the desired bound for $\partial_{x}\partial_{\mu}\mathcal{V}_{\varepsilon}$.
@@
In a similar manner to (A.2)-(A.3), applying Itô’s formulas and the convexity assumption, we obtain
$|\delta X_{1}|\leq|\phi(\xi)|\text{ a.s., }$ (A.5)
@@
$\mathbb{E}[|\delta Y_{t}|]$ $=\mathbb{E}\left[\left|\mathbb{E}\left[\partial_{x}\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1})\cdot\delta X_{1}+\tilde{\mathbb{E}}[\partial^{2}_{\mu\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1})(\tilde{X}^{t,\xi;\theta^{*}}_{1})\cdot\delta\tilde{X}_{1}]\big|\mathcal{F}_{t}\right]\right|\right]$
$\leq K_{G}\mathbb{E}[|\delta X_{1}|]+K_{G}\tilde{\mathbb{E}}[|\delta\tilde{X}_{1}|]$
$\leq 2K_{G}\mathbb{E}[|\phi(\xi)|].$ (A.6)
@@
By the standard maximum principle (or comparison principle) for parabolic equations or a straightforward application of the Feynman-Kac formula, the solution is bounded by the integral of the source term $R^{N}_{\varepsilon}$.

### v4

@@ 4 Conclusion
For optimization over $\mathbb{R}^{d}$, we establish convergence of the regularized control value to the global minimum at rate $\varepsilon\ln(1/\varepsilon)$ as $\varepsilon\to 0$. For optimization over Wasserstein space $\mathcal{P}_{2}(\mathbb{R}^{d})$, we introduce a regularized mean-field control formulation and approximate it by controlled $N$-particle systems. We prove quantitative convergence to the global optimum with rate $\frac{L_{\varepsilon}}{N}+C\varepsilon\ln\!\left(\frac{1}{\varepsilon}\right),$ as $N\to\infty$ and $\varepsilon\to 0$.
@@
Under the assumptions (a) (Regularity) and (b) (Averaged Convexity) in Proposition 3.2, the stochastic Pontryagin maximum principle with both necessary and sufficient conditions established in [11, Theorems 4.5, 4.6 & 5.1] implies that, for each $(t,\mu)\in[0,1)\times\mathcal{P}_{2}(\mathbb{R}^{d})$, $\xi\sim\mu$, the optimal control $\theta^{*}$ exists uniquely, and the optimal state process $X^{t,\xi;\theta^{*}}$ satisfies uniquely the following FBSDE:
$\begin{cases}dX^{t,\xi;\theta^{*}}_{s}=-\frac{1}{\varepsilon}Y^{t,\xi;\theta^{*}}_{s}\,ds+dW_{s},\quad X^{t,\xi;\theta^{*}}_{t}=\xi\sim\mu,\\ dY^{t,\xi;\theta^{*}}_{s}=Z^{t,\xi;\theta^{*}}_{s}\,dW_{s},\quad Y^{t,\xi;\theta^{*}}_{1}=\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1}),\\ \theta^{*}_{s}=-\frac{1}{\varepsilon}Y^{t,\xi;\theta^{*}}_{s},\quad\forall\,s\in[t,1],\end{cases}$ (A.1)
which, combined with the theories in [16, Section 5.3 and Theorem 5.7], implies that the value function $\mathcal{V}_{\varepsilon}$ is a classical solution to the master equation (1.20) satisfying Assumption 3.1 and that there holds the decoupling relation $Y^{t,\xi;\theta^{*}}_{s}=\partial_{\mu}\mathcal{V}_{\varepsilon}(s,\mathcal{L}(X^{t,\xi;\theta^{*}}_{s}))(X^{t,\xi;\theta^{*}}_{s})$ for all $s\in[t,1]$. By Lemma 3.1, the classical solution is unique.
@@
Fix an arbitrary $h\in\mathbb{R}^{d}$. Applying Itô’s formula to the scalar product $\langle\partial_{x}X^{t,x;\theta^{*}}_{s},\partial_{x}Y_{s}^{t,x;\theta^{*}}\rangle$ and taking conditional expectation, we obtain
$\langle\partial_{x}X^{t,x;\theta^{*}}_{s}h,\partial_{x}Y_{s}^{t,x;\theta^{*}}h\rangle$ $=\mathbb{E}\Big[\underbrace{\langle\partial_{x}X_{1}^{t,x;\theta^{*}}h,\,\partial_{x}\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,x,\xi;\theta^{*}}_{1})\partial_{x}X_{1}^{t,x;\theta^{*}}h\rangle}_{\geq 0\text{ by Coercivity}}+\frac{1}{\varepsilon}\int_{s}^{1}|\partial_{x}Y_{r}^{t,x;\theta^{*}}h|^{2}\,dr\big|\mathcal{F}_{s}\Big],$
@@
$d(|\partial_{x}X^{t,x;\theta^{*}}_{s}h|^{2})=2\langle\partial_{x}X_{s}^{t,x;\theta^{*}}h,d(\partial_{x}X_{s}^{t,x;\theta^{*}})h\rangle=-\frac{2}{\varepsilon}\langle\partial_{x}X_{s}^{t,x;\theta^{*}}h,\partial_{x}Y_{s}^{t,x;\theta^{*}}h\rangle\,ds\leq 0.$ (A.3)
Recall $\partial_{x}X^{t,x;\theta^{*}}_{t}=I_{d}$. Thus, it holds that $|\partial_{x}X_{s}^{t,x;\theta^{*}}|_{op}\leq|I_{d}|_{op}=1$ for all $s\in[t,1]$. Looking at $\partial_{x}Y^{t,x;\theta^{*}}$, we have
$|\partial_{x}\partial_{\mu}\mathcal{V}_{\varepsilon}(t,\mu)(x)|_{op}=|\partial_{x}Y_{t}^{t,x;\theta^{*}}|_{op}$ $=\left|\mathbb{E}\left[\partial_{x}(\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,x,\xi;\theta^{*}}_{1}))\cdot\partial_{x}X^{t,x;\theta^{*}}_{1}\big|\mathcal{F}_{t}\right]\right|_{op}$
$\leq K_{G}\mathbb{E}[|\partial_{x}X^{t,x;\theta^{*}}_{1}|_{op}|\mathcal{F}_{t}]$
$\leq K_{G},$
which yields the desired bound for $\partial_{x}\partial_{\mu}\mathcal{V}_{\varepsilon}$.
@@
Applying Itô’s formula to the scalar product $\langle\delta X_{s},\delta Y_{s}\rangle$, we have
$d\big(\langle\delta X_{s},\delta Y_{s}\rangle\big)$ $=-\frac{1}{\varepsilon}|\delta Y_{s}|^{2}\,ds+\langle\delta X_{s},\delta Z_{s}\,dW_{s}\rangle.$
To obtain an $\varepsilon$-uniform estimate, we apply the chain rule to the process $s\mapsto|\delta X_{s}|$:
$d|\delta X_{s}|=\frac{1_{|\delta X_{s}|>0}}{|\delta X_{s}|}\langle\delta X_{s},d(\delta X_{s})\rangle=-\frac{1_{|\delta X_{s}|>0}}{\varepsilon|\delta X_{s}|}\langle\delta X_{s},\delta Y_{s}\rangle\,ds,$ (A.5)
@@
$d\left(\frac{1_{|\delta X_{s}|>0}}{|\delta X_{s}|}\langle\delta X_{s},\delta Y_{s}\rangle\right)$ $=\frac{1_{|\delta X_{s}|>0}}{|\delta X_{s}|}d\big(\langle\delta X_{s},\delta Y_{s}\rangle\big)-\frac{1_{|\delta X_{s}|>0}}{|\delta X_{s}|^{2}}\langle\delta X_{s},\delta Y_{s}\rangle d|\delta X_{s}|$
$=-\frac{1_{|\delta X_{s}|>0}}{\varepsilon|\delta X_{s}|}|\delta Y_{s}|^{2}\,ds+\frac{1_{|\delta X_{s}|>0}}{|\delta X_{s}|}\langle\delta X_{s},\delta Z_{s}\,dW_{s}\rangle+\frac{1_{|\delta X_{s}|>0}}{\varepsilon|\delta X_{s}|^{3}}\langle\delta X_{s},\delta Y_{s}\rangle^{2}ds$
$\leq\frac{1_{|\delta X_{s}|>0}}{|\delta X_{s}|}\langle\delta X_{s},\delta Z_{s}\,dW_{s}\rangle,$ (A.6)
where in the last inequality we have used the Cauchy-Schwarz inequality: $|\langle\delta X_{s},\delta Y_{s}\rangle|\leq|\delta X_{s}|\cdot|\delta Y_{s}|$. Therefore, the process $\left\{\frac{1_{|\delta X_{s}|>0}}{|\delta X_{s}|}\langle\delta X_{s},\delta Y_{s}\rangle\right\}_{s\in[t,1]}$ is a supermartingale. Noticing that by the averaged convexity assumption on $G$,
$\mathbb{E}\left[\frac{1_{|\delta X_{1}|>0}}{|\delta X_{1}|}\langle\delta X_{1},\delta Y_{1}\rangle\right]$ $=\mathbb{E}\left[\frac{1_{|\delta X_{1}|>0}}{|\delta X_{1}|}\langle\delta X_{1},\partial_{x}\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1})\cdot\delta X_{1}\rangle\right]$
$\quad+\mathbb{E}\left[\frac{1_{|\delta X_{1}|>0}}{|\delta X_{1}|}\Big\langle\delta X_{1},\widetilde{\mathbb{E}}\big[\partial^{2}_{\mu\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1})(\widetilde{X}^{t,\xi;\theta^{*}}_{1})\cdot\delta\widetilde{X}_{1}\big]\Big\rangle\right]$
$\geq 0,$
we have by the supermartingale property that for all $s\in[t,1]$,
$\mathbb{E}\left[\frac{1_{|\delta X_{s}|>0}}{|\delta X_{s}|}\langle\delta X_{s},\delta Y_{s}\rangle\right]\geq 0.$
Then, integrating (A.5) from $t$ to $1$ and taking the expectations, we have
$\mathbb{E}\big[|\delta X_{1}|\big]=\mathbb{E}\big[|\phi(\xi)|\big]-\frac{1}{\varepsilon}\int_{t}^{1}\mathbb{E}\left[\frac{1_{|\delta X_{s}|>0}}{|\delta X_{s}|}\langle\delta X_{s},\delta Y_{s}\rangle\right]ds\leq\mathbb{E}\big[|\phi(\xi)|\big].$ (A.7)
@@
By the standard maximum principle (or comparison principle) for parabolic equations or a straightforward application of the Feynman-Kac formula, the solution is bounded by the source term $R^{N}_{\varepsilon}$, i.e.,